OpenStudy (anonymous):

please help me solve: 42=-3(m+6)+m

OpenStudy (mathmale):

In your shoes I'd begin by distributing that -3 over (m+6): multiply each m and 6 by -3. Please type out the resulting equation.

OpenStudy (anonymous):

-3m-18

OpenStudy (mathmale):

Good, and so your 42=-3(m+6)+m then becomes 42=-3m-18 +m.

OpenStudy (mathmale):

Please add the "like terms" on the right. In other words, combine -3m and +m.

OpenStudy (anonymous):

4m

OpenStudy (mathmale):

I see -3m + 1m, or: -3m +1m ----- ?

OpenStudy (anonymous):

2m

OpenStudy (mathmale):

The -3m has a larger coefficient (ignoring the sign) than does +1m. Therefore, the combination of -3m and +1m results in -2m. Can you convince yourself of this?

OpenStudy (anonymous):

oh ok so -2

OpenStudy (mathmale):

Then your 42=-3(m+6)+m becomes 42=-2m -18. What do you think is the next step that will help you solve this equation for m?

OpenStudy (mathmale):

Solve for m: 42=-2m -18.

OpenStudy (anonymous):

-18-24